# Distributed Differentially Private Data Analytics via Secure Sketching

### v1

@@ Appendix B Privacy Proofs
Otherwise, $\mathcal{M}(\mathbf{A})=T_{\mathbf{S}}(R_{\sigma}(\mathbf{a}_{1}),\dots,R_{\sigma}(\mathbf{a}_{n}))=\mathbf{S}(\mathbf{A}+\mathbf{G})$ where $\mathbf{a}_{i}$ denotes row $i$ of $\mathbf{A}\in\mathbb{R}^{n\times d}$ and $\mathbf{G}\leftarrow\mathcal{N}(0,\sigma^{2})^{n\times d}$, is $(\varepsilon,\delta)$ differentially private.
Consider first $d=1$ and let mechanism $\mathcal{M}_{1}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be defined as $\mathcal{M}_{1}(\mathbf{a})=\mathbf{S}(\mathbf{a}+\mathbf{g})$ for $\mathbf{a}\in\mathbb{R}^{n}$, with $\mathbf{g}\leftarrow\mathcal{N}(0,\sigma^{2})^{n}$.
This is because $\mathbf{S}$ is sampled such that it has only one non-zero entry per column; thus, the columns of $\mathbf{S}$ are orthogonal. This means that the entries of $(\mathbf{S}(\mathbf{a}+\mathbf{g}))_{j}$ and $(\mathbf{S}(\mathbf{a}+\mathbf{g}))_{j^{\prime}}$ for $j\neq j^{\prime}$, have disjoint support. Therefore, the privacy guarantee can be analyzed independently for each entry of $\mathbf{S}(\mathbf{a}+\mathbf{g})$. Then the overall privacy guarantee for $\mathcal{M}_{1}$ is the guarantee of $\mathbf{S}_{i}(\mathbf{a}+\mathbf{g})$, where $i$ is the index of the row in $\mathbf{S}$ with the fewest non-zero entries (i.e. the one where the least noise is added from $\mathbf{g}$).
Notice that the $\ell_{2}$-sensitivity of algorithm $T_{\mathbf{S}}$ is $\Delta_{2}T_{\mathbf{S}}=\max_{\mathbf{A},\mathbf{A}^{\prime}}\lVert\mathbf{S}\mathbf{A}-\mathbf{S}\mathbf{A}^{\prime}\rVert_{2}=2\eta\sqrt{m}$, where $\mathbf{A},\mathbf{A}^{\prime}\in\mathbb{R}^{n\times d}$ differ in a single row, and every entry of $\mathbf{A},\mathbf{A}^{\prime}$ is bounded above by $\eta$.
If $\delta/d=\delta>m\exp(-(n-t^{\prime})/8m)$, we sample the individual entries of $\mathbf{g}$ from $\mathcal{N}(0,\sigma^{2})$ with
$\sigma^{2}=\frac{1}{n-t^{\prime}}\cdot\frac{4m\eta^{2}\ln(1.25/(\delta-m\exp(-(n-t^{\prime})/8m)))m}{\varepsilon^{2}}.$
Denote by $E$ the event that the support of $\mathbf{S}(a+g)_{j}$ has at least $\frac{n-t^{\prime}}{2m}$ uncorrupted clients, whose set of indices we denote $N$. Setting $\gamma=1/2$, Lemma B.1 gives that $E$ occurs with probability at least $1-m\exp(-\frac{n-t^{\prime}}{8m})$. In event $E$, due to the stability of Gaussians, we can formulate $\mathbf{S}(\mathbf{a}+\mathbf{g})_{j}=(\mathbf{S}\mathbf{a})_{j}+(\mathbf{S}\mathbf{g})_{j}$. Due to symmetry of Gaussians around the origin, $(\mathbf{S}\mathbf{g})_{j}$ is Gaussian distributed with mean zero and variance:
$\tilde{\sigma}^{2}$ $\geq\frac{n-t^{\prime}}{2m}\sigma^{2}$
$=\frac{n-t^{\prime}}{2m}\cdot\frac{4m\eta^{2}\ln(1.25/(\delta-m\exp(-(n-t^{\prime})/8m)))m}{\varepsilon^{2}(n-t^{\prime})}$
$=\frac{2m\eta^{2}\ln(1.25/(\delta-m\exp(-(n-t^{\prime})/8m)))}{\varepsilon^{2}}.$
Then in event $E$, $\mathcal{M}_{1}$ is $(\varepsilon,\delta-m\exp(-(n-t^{\prime})/8m)))$ differentially private using the Gaussian mechanism (Lemma 2.3). Since event $E$ does not occur with probability at most $m\exp(-(n-t^{\prime})/8m)$, $\mathcal{M}_{1}$ is $(\varepsilon,\delta-m\exp(-(n-t^{\prime})/8m)))$ differentially private except with probability $m\exp(-(n-t^{\prime})/8m)$. Then $\mathcal{M}_{1}$ is $(\varepsilon,\delta)$-differentially private.
@@
We now generalize to arbitrary choices of $d$. Note that now $\mathbf{S}_{i}(\mathbf{A}+\mathbf{G})$ is a d-dimensional vector. Sequential Composition (Lemma C.1) then gives that $\mathcal{M}$ is $(\varepsilon,\delta)$ differentially private, if we sample $\mathbf{G}$’s entries from $\mathcal{N}(0,\sigma^{2})$ with
$\sigma^{2}=\frac{4\eta^{2}\ln(1.25/(\delta/d-m\exp(-(n-t^{\prime})/8m)))m^{2}d^{2}}{\varepsilon^{2}(n-t^{\prime})}$
Proof of Theorem 3.4.
@@
Let $\varepsilon$, $\delta$, $R_{\sigma}$ and $T_{\mathbf{S}}$ be given as described in the theorem. We then need to prove that mechanism $\mathcal{M}:{\mathbb{R}}^{n\times d}\to{\mathbb{R}}^{m\times d}$ defined by $\mathcal{M}(\mathbf{A})=T_{\mathbf{S}}(R_{\sigma}(\mathbf{a}_{1}),\dots,R_{\sigma}(\mathbf{a}_{n}))$ is $(\varepsilon,\delta)$ differentially private. If $\delta/d=\delta\leq m\exp(-(n-t^{\prime})/8m)$, then the $R_{\sigma}$ outputs $0^{d}$, which is entirely independent of the input and is therefore $(\varepsilon,\delta)$ differentially private.
@@
Otherwise, $\mathcal{M}(\mathbf{A})=T_{\mathbf{S}}(R_{\sigma}(\mathbf{a}_{1}),\dots,R_{\sigma}(\mathbf{a}_{n}))=\frac{1}{\sqrt{s}}\sum_{i\in[s]}\mathbf{S}_{i}(\mathbf{A}+\mathbf{G}_{i})$ where $\mathbf{a}_{i}$ denotes row $i$ of $\mathbf{A}\in\mathbb{R}^{n\times d}$ and $\mathbf{G}_{i}\leftarrow\mathcal{N}(0,\sigma^{2})^{n\times d}$. Here, $\mathbf{S}$ has been decomposed such that $\mathbf{S}=\frac{1}{\sqrt{s}}\sum_{i}^{s}\mathbf{S}_{i}$, where $\mathbf{S}_{i}\in\{-1,0,1\}^{m\times n}$ is a matrix with only one non-zero entry per column.
@@
We argue $(\varepsilon/s,\delta/s)$-DP based on Theorem B.2. Note that the non-zero entry per column in each $\mathbf{S}_{i}$ is not sampled uniformly at random from all possible rows; rather, for $\mathbf{S}_{i}$ is sampled uniformly at random from the remaining $n-i$ rows (removing the one that was already chosen to be non-zero, sampling without replacement). Therefore, we replace $n$ with $n-s$ [^3] Since $n\gg m>s$, subtracting $s$ will not have a large effect. when applying Theorem B.2.
@@
Appendix C Utility Proofs
Lemma C.1 (Sequential Composition [23]).
Lemma C.2.

### v2

@@ 4.2. Privacy Preserving Mechanism for Oblivious Johnson Lindenstrauss Transforms
Otherwise, $\mathcal{M}(\mathbf{A})=T_{\mathbf{S}}(R_{D}(\mathbf{a}_{1}),\dots,R_{D}(\mathbf{a}_{n}))=\mathbf{S}(\mathbf{A}+\mathbf{G})$ where $\mathbf{a}_{i}$ denotes row $i$ of $\mathbf{A}\in\mathbb{R}^{n\times d}$ and $\mathbf{G}\leftarrow D^{n\times d}$, and we now prove that this is $(\varepsilon,\delta)$ differentially private.
@@
Note that $\mathbf{S}_{i}(\mathbf{A}+\mathbf{G})$ is a d-dimensional vector, when $\mathbf{S}_{i}$ refers to the $i$-th row of $\mathbf{S}$. If $(\mathbf{S}_{i}(\mathbf{A}+\mathbf{G}))_{j}$ is $(\varepsilon/d,\delta/d)$-DP, then sequential composition (Lemma 3.5) gives that $\mathcal{M}$ is $(\varepsilon,\delta)$ differentially private.
Now consider $d=1$ and let mechanism $\mathcal{M}_{1}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be defined as $\mathcal{M}_{1}(\mathbf{a})=\mathbf{S}(\mathbf{a}+\mathbf{g})$ for $\mathbf{a}\in\mathbb{R}^{n}$, with $\mathbf{g}\leftarrow D^{n}$.
$\mathbf{S}$ is sampled such that it has only one non-zero entry per column; thus, the columns of $\mathbf{S}$ are orthogonal. This means that the entries of $(\mathbf{S}(\mathbf{a}+\mathbf{g}))_{j}$ and $(\mathbf{S}(\mathbf{a}+\mathbf{g}))_{j^{\prime}}$ for $j\neq j^{\prime}$, have disjoint support. Therefore, the privacy guarantee can be analyzed independently for each entry of $\mathbf{S}(\mathbf{a}+\mathbf{g})$.
Denote by $E$ the event that the support of $\mathbf{S}(a+g)_{j}$ has at least $\frac{n-t^{\prime}}{2m}$ uncorrupted clients, whose set of indices we denote $N$. We can formulate $\mathbf{S}(\mathbf{a}+\mathbf{g})_{j}=(\mathbf{S}\mathbf{a})_{j}+(\mathbf{S}\mathbf{g})_{j}=(\mathbf{S}\mathbf{a})_{j}+\sum_{k\in[n]}\mathbf{S}_{j,k}\mathbf{g}_{k}=(\mathbf{S}\mathbf{a})_{j}+\sum_{k\in N}\mathbf{S}_{j,k}\mathbf{g}_{k}+\sum_{k\in[n]\setminus N}\mathbf{S}_{j,k}\mathbf{g}_{k}$. Also note that $\sum_{k\in N}\mathbf{S}_{j,k}\mathbf{g}_{k}$ contains at least $\frac{n-t^{\prime}}{2m}$ terms in event $E$, the sum of which suffice for guaranteeing DP.
Due to the fact that $D$ is symmetric, $-1\cdot\mathbf{g}_{k}$ follows the same distribution as $\mathbf{g}_{k}$, and $\sum_{k\in N}\mathbf{S}_{j,k}\mathbf{g}_{k}$ follows the same distribution as $\sum_{k\in N}\mathbf{g}_{k}$. Due to the infinite divisibility of $D^{\prime}$, the sum of the first $\frac{n-t^{\prime}}{2m}$ terms of $\sum_{k\in N}\mathbf{g}_{k}$ are samples from $D^{\prime}$.
Then in event $E$, $\mathcal{M}_{1}$ is $(\varepsilon,\delta-m\exp(-(n-t^{\prime})/8m)))$ differentially private using the additive noise mechanism that adds noise from $D^{\prime}$. Setting $\gamma=1/2$, Lemma 4.5 gives that $E$ occurs with probability at least $1-m\exp(-\frac{n-t^{\prime}}{8m})$. Since event $E$ does not occur with probability at most $m\exp(-(n-t^{\prime})/8m)$, $\mathcal{M}_{1}$ is $(\varepsilon,\delta-m\exp(-(n-t^{\prime})/8m)))$ differentially private except with probability $m\exp(-(n-t^{\prime})/8m)$. Then $\mathcal{M}_{1}$ is $(\varepsilon,\delta)$-differentially private.
@@
Let $\varepsilon$, $\delta$, $R_{D}$ and $T_{\mathbf{S}}$ be given as described in the theorem. If $\delta/d\leq m\exp(-(n-s-t^{\prime})/8m)$, then the $R_{D}$ outputs $0^{d}$, which is entirely independent of the input and is therefore $(\varepsilon,\delta)$ differentially private.
@@
Otherwise, $\mathcal{M}(\mathbf{A})=T_{\mathbf{S}}(R_{D}(\mathbf{a}_{1}),\dots,R_{D}(\mathbf{a}_{n}))=\frac{1}{\sqrt{s}}\sum_{i\in[s]}\mathbf{S}_{i}(\mathbf{A}+\mathbf{G}_{i})$ where $\mathbf{a}_{i}$ denotes row $i$ of $\mathbf{A}\in\mathbb{R}^{n\times d}$ and $\mathbf{G}_{i}\leftarrow D^{n\times d}$. Here, $\mathbf{S}$ has been decomposed such that $\mathbf{S}=\frac{1}{\sqrt{s}}\sum_{i}^{s}\mathbf{S}_{i}$, where
$\mathbf{S}_{i}\in\{-1,0,1\}^{m\times n}$ is a matrix with only one non-zero entry per column.
@@
We argue $(\varepsilon/s,\delta/s)$-DP based on Lemma 4.6. Note that the non-zero entry per column in each $\mathbf{S}_{i}$ is not sampled uniformly at random from all possible rows; rather, for $\mathbf{S}_{i}$ is sampled uniformly at random from the remaining $n-i$ rows (removing the one that was already chosen to be non-zero, sampling without replacement). Therefore, we replace $n$ with $n-s$ when applying Lemma 4.6.
@@
One additive mechanism to which this can easily be applied is the Gaussian mechanism. To do so, $D$ is a normal distribution with variance described in the following corollary.
Corollary 4.7.
Let $\varepsilon\geq 0$, $\delta\in(0,1)$, $t^{\prime}<n$, $m\in[n]$, $s\in[n]$ and $D$ be the Gaussian distribution $\mathcal{N}(0,\sigma^{2})$ with:
$\sigma^{2}=\frac{4s^{3}\eta^{2}\ln(1.25s/(\delta/d-m\exp(\frac{-(n-s-t^{\prime})}{8m})))md^{2}}{\varepsilon^{2}(n-s-t^{\prime})}$
Let $\mathbf{S}\sim\mathcal{D}_{m,n,s}^{\mathtt{sketch}}$. Then as long as input values are bounded above by $\eta$, tuple of algorithms $(R_{D},T_{\mathbf{S}},A)$, as defined by equations 2 and 3 and for any $A$, is $(\varepsilon,\delta)$-differentially private in the trusted computation model for differential privacy with $t^{\prime}$ corrupt clients.
Proof.
We apply Theorem 4.4, setting $D$ to a Gaussian distribution with the variance specified in the corollary. First note that zero-centered Gaussians are symmetric around the origin. Due to infinite divisibility of Gaussians, $\sum_{i\in[(n-s-t^{\prime})/(2m)]}g_{i}=G$ when $\mathbf{g}_{i}\sim D$ is Gaussian distributed with variance
$\sum_{i\in[(n-s-t^{\prime})/(2m)]}\sigma^{2}=(n-s-t^{\prime})\sigma^{2}/(2m)$
Notice that the $\ell_{2}$-sensitivity of algorithm $T_{\mathbf{S}}$ is $\Delta_{2}T_{\mathbf{S}}=\max_{\mathbf{A},\mathbf{A}^{\prime}}\lVert\mathbf{S}\mathbf{A}-\mathbf{S}\mathbf{A}^{\prime}\rVert_{2}=2\eta\sqrt{s}$, where $\mathbf{A},\mathbf{A}^{\prime}\in\mathbb{R}^{n\times d}$ differ in a single row, and every entry of $\mathbf{A},\mathbf{A}^{\prime}$ is bounded above by $\eta$.
Based on the Gaussian mechanism (Lemma 3.3), the additive noise mechanism adding Gaussian noise satisfies $(\varepsilon/(sd),\delta/(sd)-m\exp(-(n-s-t^{\prime})/8m)))$-DP for function $f(\mathbf{A})=(\mathbf{S}\mathbf{A})_{j}+\mathbf{n}$ with $\mathbf{n}\sim\text{Lap}(0,2\eta m^{2}d/\varepsilon)^{m}$ on input $\mathbf{A}\in\mathbb{R}^{n\times d}$, where the subscript refers to one column of $\mathbf{S}\mathbf{A}$.
If we would like to obtain a result for mechanisms with $\delta=0$, we can instead use $R^{\text{dense}}_{D}$ and use a dense sketching matrix $\mathbf{S}\in\{-1,1\}^{m\times n}$.
Theorem 4.8.
Let $\varepsilon\geq 0$, $\delta\in[0,1)$, $m\in[n]$, $t^{\prime}<n$, $s\in[m]$. Let $\mathbf{S}\sim\{-1,1\}^{m\times n}$ be a dense, uniformly sampled matrix. Also let $R$ add noise $g_{i}$ sampled from a symmetric distribution $D$ such that $\sum_{i\in[n/m-t^{\prime}]}g_{i}=G$. $G$ follows a distribution $D^{\prime}$ that defines an additive noise mechanism that satisfies $(\varepsilon/(md),\delta/(md))$-DP for function $f(\mathbf{A})=(\mathbf{S}\mathbf{A})_{j}+\mathbf{n}$ for $\mathbf{n}\sim D^{\prime m}$ on input $\mathbf{A}\in\mathbb{R}^{n\times d}$, where the subscript refers to one column of $\mathbf{S}\mathbf{A}$. Then as long as input values are bounded above by $\eta$, dense sketching mechanism $(R_{D}^{\text{dense}},T_{\mathbf{S}}^{\text{dense}},A)$, as defined by equations 1 and 3 and for any $A$, is $(\varepsilon,\delta)$-differentially private in the LTM for differential privacy with $t^{\prime}$ corrupt clients.
Proof.
We again begin by proving a simplified version of the theorem first.
Lemma 4.9.
Let $\varepsilon\geq 0$, $\delta\in[0,1)$, $m\in[n]$, $t^{\prime}<n$. Let $\mathbf{S}\in\{-1,0,1\}^{m\times n}$ with one non-zero entry per column and $n/m$ non-zero entries per row. Also let noise $g_{i}$ be sampled from a symmetric distribution $D$ such that $\sum_{i\in[(n)/m-t^{\prime}]}g_{i}=G$. $G$ follows a distribution $D^{\prime}$ that defines an additive noise mechanism that satisfies $(\varepsilon/d,\delta/d)$-DP for function $f(\mathbf{A})=(\mathbf{S}\mathbf{A})_{j}+\mathbf{n}$ with $\mathbf{n}\sim D^{\prime m}$ on input $\mathbf{A}\in\mathbb{R}^{n\times d}$, where the subscript refers to one column of $\mathbf{S}\mathbf{A}$. Then as long as input values are bounded above by $\eta$, tuple of algorithms $(R^{\text{dense}}_{D},T_{\mathbf{S}},A)$, as defined by equations 1 and 3 and for any $A$, is $(\varepsilon,\delta)$-differentially private in the trusted computation model for differential privacy with $t^{\prime}$.
Proof of Lemma 4.9.
Let $\varepsilon$, $\delta$, $R^{\text{dense}}_{D}$ and $T_{\mathbf{S}}$ be given as described in the theorem.
We show that $\mathcal{M}(\mathbf{A})=T_{\mathbf{S}}(R(\mathbf{a}_{1}),\dots,R(\mathbf{a}_{n}))=\mathbf{S}(\mathbf{A}+\mathbf{G})$ where $\mathbf{a}_{i}$ denotes row $i$ of $\mathbf{A}\in\mathbb{R}^{n\times d}$ and $\mathbf{G}\leftarrow D^{n\times d}$, is $(\varepsilon,\delta)$ differentially private.
Note that $\mathbf{S}_{i}(\mathbf{A}+\mathbf{G})$ is a d-dimensional vector, when $\mathbf{S}_{i}$ refers to the $i$-th row of $\mathbf{S}$. If $(\mathbf{S}_{i}(\mathbf{A}+\mathbf{G}))_{j}$ is $(\varepsilon/d,\delta/d)$-DP, then sequential composition (Lemma 3.5) gives that $\mathcal{M}$ is $(\varepsilon,\delta)$ differentially private.
Consider $d=1$ and let mechanism $\mathcal{M}_{1}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be defined as $\mathcal{M}_{1}(\mathbf{a})=\mathbf{S}(\mathbf{a}+\mathbf{g})$ for $\mathbf{a}\in\mathbb{R}^{n}$, with $\mathbf{g}\leftarrow D^{n}$.
$\mathbf{S}$ is sampled such that it has only one non-zero entry per column; thus, the columns of $\mathbf{S}$ are orthogonal. This means that the entries of $(\mathbf{S}(\mathbf{a}+\mathbf{g}))_{j}$ and $(\mathbf{S}(\mathbf{a}+\mathbf{g}))_{j^{\prime}}$ for $j\neq j^{\prime}$, have disjoint support. Therefore, the privacy guarantee can be analyzed independently for each entry of $\mathbf{S}(\mathbf{a}+\mathbf{g})$.
Note that the support of $\mathbf{S}(a+g)_{j}$ has at least $\frac{n}{m}-t^{\prime}$ uncorrupted clients, whose set of indices we denote $N$. We can formulate $\mathbf{S}(\mathbf{a}+\mathbf{g})_{j}=(\mathbf{S}\mathbf{a})_{j}+(\mathbf{S}\mathbf{g})_{j}=(\mathbf{S}\mathbf{a})_{j}+\sum_{k\in[n]}\mathbf{S}_{j,k}\mathbf{g}_{k}=(\mathbf{S}\mathbf{a})_{j}+\sum_{k\in N}\mathbf{S}_{j,k}\mathbf{g}_{k}+\sum_{k\in[n]\setminus N}\mathbf{S}_{j,k}\mathbf{g}_{k}$.
Due to the fact that $D$ is symmetric, $-1\cdot\mathbf{g}_{k}$ follows the same distribution as $\mathbf{g}_{k}$, and $\sum_{k\in N}\mathbf{S}_{j,k}\mathbf{g}_{k}$ follows the same distribution as $\sum_{k\in N}\mathbf{g}_{k}$. Due to the infinite divisibility of $D^{\prime}$, the sum of the first $\frac{n}{m}-t^{\prime}$ terms of $\sum_{k\in N}\mathbf{g}_{k}$ are samples from $D^{\prime}$.
Then $\mathcal{M}_{1}$ is $(\varepsilon,\delta)$-differentially private. ∎
Let $\varepsilon$, $\delta$, $R_{D}^{\text{dense}}$ and $T_{\mathbf{S}}^{\text{dense}}$ be given as described in the theorem.
We know $\mathcal{M}(\mathbf{A})=T_{\mathbf{S}}(R_{D}(\mathbf{a}_{1}),\dots,R_{D}(\mathbf{a}_{n}))=\frac{1}{\sqrt{m}}\sum_{i\in[m]}\mathbf{S}_{i}(\mathbf{A}+\mathbf{G}_{i})$ where $\mathbf{G}_{i}\leftarrow D^{n\times d}$. Recall from Equation 3 that $\mathbf{S}$ has been decomposed such that $\mathbf{S}=\frac{1}{\sqrt{m}}\sum_{i}^{m}\mathbf{S}_{i}$, where $\mathbf{S}_{i}\in\{-1,0,1\}^{m\times n}$ is a matrix with only one non-zero entry per column and $n/m$ non-zero entries per row.
We argue $(\varepsilon/s,\delta/s)$-DP based on Lemma 4.9.
To instantiate this second theorem, $D$ can be chosen as the distribution formulated as the difference of two samples of the gamma distribution.
Corollary 4.10.
Let $\varepsilon\geq 0$, $t^{\prime}<n$, $m\in[n]$, $s\in[n]$, and let $D$ be the distribution formulated as the difference of two samples of gamma distribution $\Gamma(\frac{1}{n/m-t^{\prime}},b)$ with parameter
$b=2\eta m^{2}d/\varepsilon$
Let $\mathbf{S}\sim\{-1,1\}^{m\times n}$ sampled uniformly. Then as long as input values are bounded above by $\eta$, $(R_{D}^{\text{dense}},T_{\mathbf{S}},A)$ is $\varepsilon$-differentially private in the trusted computation model for differential privacy with $t^{\prime}$ corrupt clients $C^{\prime}$.
Proof.
We apply Theorem 4.8, setting $D$ to the difference between two samples of the gamma distribution $\Gamma(\frac{1}{n/m-t^{\prime}},2\eta m^{2}d/\varepsilon)$. Note that the resulting distribution is centered at 0 and symmetric. Due to infinite divisibility of Laplace distributions, $\sum_{i\in[n/m-t^{\prime}]}g_{i}=G$ when $\mathbf{g}_{i}\sim D$ is Laplace distributed with mean zero and scale parameter $b=2\eta m^{2}d/\varepsilon$.
Notice that the $\ell_{1}$-sensitivity of algorithm $T_{\mathbf{S}}$ is
$\Delta_{1}T_{\mathbf{S}}=\max_{\mathbf{A},\mathbf{A}^{\prime}}\lVert\mathbf{S}\mathbf{A}_{j}-\mathbf{S}\mathbf{A}^{\prime}_{j}\rVert_{1}=2\eta m$, where $\mathbf{A},\mathbf{A}^{\prime}\in\mathbb{R}^{n\times d}$ differ in a single row, and every entry of $\mathbf{A},\mathbf{A}^{\prime}$ is bounded above by $\eta$.
Based on the Laplace mechanism (Lemma 3.4), the additive noise mechanism adding Laplace noise satisfies $\varepsilon/(dm)$-DP for function $f(\mathbf{A})=(\mathbf{S}\mathbf{A})_{j}+\mathbf{n}$ with $\mathbf{n}\sim\text{Lap}(0,2\eta m^{2}d/\varepsilon)^{m}$ on input $\mathbf{A}\in\mathbb{R}^{n\times d}$, where the subscript refers to one column of $\mathbf{S}\mathbf{A}$.
Corollary 4.11 below follows directly from Theorem 4.4 and Lemma 4.3.
Corollary 4.11.
Let $(R_{D},T_{\mathbf{S}},A)$ be the tuple of algorithms above. If $T_{\mathbf{S}}$ is computed correctly using an MPC protocol $\Pi_{\mathbf{S}}$ with perfect security tolerating $t$ semi-honest corruptions, then the tuple of algorithms $(R_{D},\Pi_{\mathbf{S}},A)$ is $(\varepsilon,\delta)$-differentially private with $t^{\prime}$ corrupt clients.
4.3. Cryptographic Assumptions and Relations to the Shuffle Model
The LTM can be securely distributed using simple cryptographic techniques for secure multiparty computation (MPC), such as linear secret-sharing (LSS). MPC ensures that no adversary corrupting all but one server learns more than the computation’s output. (Dwork et al., 2006a) first explored combining differential privacy and MPC, focusing on distributed noise generation requiring server interaction. (Cheu and Yan, 2023) establish lower bounds for non-interactive multi-server mechanisms, while (Damgård et al., 2023) propose an interactive MPC protocol for selection in distributed trust models.
The linear transforms can be non-interactively computed locally by servers and are also known to all participating parties, that is, once initiated, the output of the protocol is fully deterministic. Linear transforms can be computed non-interactively by servers and are publicly known, making the protocol fully deterministic once initiated. This enables implementing the LTM with multiple central servers under the minimal assumption that at least one server remains honest. Using LSS has key advantages: clients send a single message per server proportional to the input size and sparsity $s$ of the transformation matrix $\mathbf{S}$, servers communicate only with the data analyst, and the total workload scales only with the number of servers and matrix sparsity. Since LSS supports information-theoretic security, our system remains secure even against quantum adversaries.
We require this secret sharing scheme to be information-theoretically secure. We use an additive secret sharing scheme, such that clients $C_{1},\dots,C_{n}$ generate $k$ shares of their noisy inputs $x^{1},\dots,x^{n}$, where $k$ is the number of servers that compute the linear transformation. We use $[x^{i}]$ to denote a secret sharing of some client $C_{i}$’s noisy input $x^{i}$ for some $i\in[n]$. For some field $\mathbb{F}$ of size $p$ and some prime number $p$, $[x^{i}]$ consists of shares $x^{i}_{1},\dots,x^{i}_{k}\in\mathbb{F}$ such that $\sum_{j=1}^{k}x_{j}^{i}=x^{i}$. To split a secret into $k$ shares, a client can sample $k-1$ random field elements $x^{i}_{1},\dots,x^{i}_{k-1}$ and compute the last secret share as $x^{i}-\sum_{j=1}^{k-1}x_{j}^{i}$. Therefore, it also seems intuitive that an adversary that sees all but one of the shares knows nothing about the input. For every $j\in[k]$, server $S_{j}$ receives $\{x_{j}^{i}\}_{i\in n}$ from the respective parties. We define these shares such that the security of our distributed protocol does not rely on any computational assumption.
Since the linear transformation is public, servers can apply the transformation locally to their secret shares without any need to communicate with each other. Each server then reveals their shares of the resulting linear transformation, such that the linear sketch can be revealed in the clear. For more details on MPC based on secret sharing, see e.g., (Cramer et al., 2015).
Secure aggregation and the shuffle model are two instantiations of intermediate trust models for differential privacy. Secure aggregation (Goryczka and Xiong, 2017; Bonawitz et al., 2017; Mugunthan et al., 2019; Apple and Google, 2021; Talwar et al., 2023) is a special case of the LTM, where the linear transformation applied is a sum, which is useful in federated learning. The shuffle model (Bittau et al., 2017; Erlingsson et al., 2019; Cheu et al., 2019) can be distributed among multiple servers using cryptographic protocols such as mixnets.
However, implementing shuffles is more complex than linear transformations, requiring servers to communicate over a chain (increasing latency), computationally intensive public-key cryptography, which in turns requires computational assumptions. In contrast, LSS enables efficient, unconditional security and deterministic computation. [^1] The linear transformations we use are random but public and sampled before the computation.
5. Numerical Linear Algebra in LTM
In this section we give the utility guarantees of our mechanism. We start with giving generic distortion bounds relating the spectrum of the noisy matrix to the spectrum of the original non-private matrix. Parameterizations of this mechanism depend on the underlying class of subspace embeddings. In this utility analysis, we limit outselves to OSNAPs with $m$ and $s$ chosen according to (Cohen, 2016); other trade-offs are possible and can be found in Section 3. Applications to specific problems such as regression and low rank approximation are given towards the end of the section.
Here we present the formal utility statements for ridge regression and low rank approximation. We begin by introducing a useful helper lemma.
Lemma 5.1.